# Multi-factor modeling of chlorophyll-a in South China’s subtropical reservoirs using long-term monitoring data for quantitative analysis

### v2

@@ Abstract
Eutrophication and harmful algal blooms, driven by complex interactions among nutrients and climate, threaten freshwater ecosystems globally, particularly in densely populated Asian regions. Understanding interactions among water temperature, nutrient levels, and chlorophyll-a (Chl-a) dynamics is crucial for addressing eutrophication in freshwater ecosystems. However, many existing studies tend to oversimplify these relationships, often neglecting the non-linear effects and long-term temporal variations. Here, we conducted multi-year field monitoring (2020-2024) of key environmental factors, including total nitrogen (TN), total phosphorus (TP), water temperature, and Chl-a, across three reservoirs in Guangdong Province, China: Tiantangshan (S1), Baisha River (S2), and Meizhou (S3). Chl-a concentrations showed significant spatiotemporal variability, ranging from 1.2 to 11.8 $\mu\text{g}/\text{L}$, with a general increasing trend indicative of progressing eutrophication. Strong positive correlations were found between Chl-a and TN, TP, and temperature. Long-term data revealed TN as a more influential driver than TP for Chl-a proliferation in these systems. Based on the collected data, we developed and calibrated a dynamic multi-factor hydro-ecological model. The model accurately reproduced the observed Chl-a patterns (R2 > 0.85), identifying synergistic effects between temperature and nutrients. The model offers a robust theoretical basis for predicting Chl-a dynamics and supports science-informed management strategies to mitigate eutrophication in subtropical reservoirs.
@@
3Ocean College, Jiangsu University of Science and Technology, Zhenjiang 212100, PR China
@@ 1. Introduction
Aquatic ecosystems play a fundamental role in supporting life on Earth, with their structure and functioning shaped by complex interactions among key environmental factors, particularly water temperature, nutrient availability (e.g., total nitrogen (TN) and total phosphorus (TP)), and algal biomass, typically represented by chlorophyll-a (Chl-a). In recent decades, global climate change and intensified human activity have profoundly altered these ecosystems, driving more frequent extreme hydrological events, accelerating eutrophication, and contributing to widespread biodiversity loss (Fong et al., 2025; Priya et al.,, 2023; Hader and Barnes, 2019; Menden-Deuer et al., 2025; Soanar et al., 2024; Zhao et al., 2025; Hirata et al., 2025; Boyacioglu et al., 2024). In Asia, rapid urbanization and agricultural intensification have exacerbated nutrient pollution in freshwater bodies. Studies from South Asia, including Pakistan and Bangladesh, document severe eutrophication pressures in river basins and lakes, highlighting a regional crisis that demands locally calibrated understanding and solutions (Haq and Muhammad., 2023; Muhammad., 2023; Tokatlı et al., 2025, 2023; Muhammad et al., 2024). The situation in China’s subtropical regions is similarly pressing, with reservoirs experiencing frequent algal blooms due to combined nutrient loading and warming trends.
@@
Although multi-factor modeling in aquatic ecology has advanced in recent years, many models remain limited in predictive and explanatory power. Early studies typically focused on linear, single-factor relationships, such as nutrient concentrations and chlorophyll-a growth. However, the emerging field of complexity science has prompted a shift toward integrative approaches that capture interactions among multiple drivers. For instance, Liu et al. (Liu et al., 2021) employed a coupled hydrodynamic-ecological model to quantify the interactive effects of TN, TP, total suspended solids (TSS), and light availability on phytoplankton competition in Chagan Lake, identifying TSS as a primary driver of cyanobacterial succession. Similarly, Qian et al. (Qian et al., 2024) applied deep learning to model algal bloom dynamics in Taihu Lake, revealing synergistic effects between thermal stratification and nutrient input. Zhang et. al. (Zhang et al., 2024) explored spatial patterns of TP and chlorophyll-a in reservoirs, highlighting the influence of geographic and meteorological factors such as latitude, slope, and temperature on chlorophyll-a variability. Remote sensing studies in subtropical reservoirs, such as the Nandoni reservoir in South Africa, have further corroborated the utility of high-resolution monitoring in capturing the spatial heterogeneity of Chl-a, often revealing higher concentrations near inflows and reservoir edges, which aligns with the need for spatially explicit understanding of eutrophication drivers (Mpakairi et al., 2024). Systematic reviews emphasize the growing effort to develop integrated models that couple watershed processes with in-lake dynamics to support management (Shi et al., 2024; Buta et al., 2023). In South Asia, research in the Ganges-Brahmaputra and Indus River basins has also highlighted the critical role of anthropogenic nutrient loads and climate variability in driving aquatic ecosystem changes (Muhammad et al., 2024; Tokatlı et al., 2023).
@@
While these studies contribute valuable data and demonstrate the potential of advanced modeling techniques, they often rely heavily on statistical or machine learning approaches, offering limited insight into the underlying ecological mechanisms  (Liu et al., 2021; Qian et al., 2024; Zhang et al., 2024). Furthermore, many models are calibrated with short-term data and lack the validation against long-term observations needed to assess their performance under varying climatic conditions and episodic events, such as extreme rainfall which is common in subtropical monsoonal regions (Buta et al., 2023). Moreover, traditional models such as the Water Erosion Prediction Project (WEPP), although effective in simulating physical processes like soil erosion, fall short in capturing hydro-ecological interactions (Flanagan et al., 2012; Chen et al., 2025). Some progress has been made by integrating robust optimization and scenario analysis into water resource management, for example, Xu et al. (Xu et al., 2022) incorporated multi-objective optimization and copula-based uncertainty quantification into cascade reservoir planning, but the ecological models remain relatively simplified. In contrast, emerging dynamic multi-factor models offer promising tools to explore non-linear feedbacks and temporal continuity among interacting ecological variables, including chlorophyll-a growth, nutrient cycling, and temperature dynamics. Despite their potential, these models are still in early stages of development and are seldom applied to real-world aquatic systems.
@@
To address these gaps, this study has two clear objectives: (1) to quantify the individual and interactive effects of water temperature, TN, and TP on Chl-a dynamics in three contrasting subtropical reservoirs using a four-year monitoring dataset; and (2) to develop, calibrate, and validate a process-based dynamic model that can simulate these interactions and be used for scenario analysis to support management. We propose a dynamic, multi-factor modeling approach to investigate the interactions among water temperature, TN, TP, and chlorophyll-a in three freshwater reservoirs in Guangdong, China. By leveraging long-term monitoring data and integrating ecological mechanisms into the modeling framework, this work aims to deepen our understanding of eutrophication processes and provide a scientific basis for informed aquatic ecosystem management.
@@ 2.1. Research sites and sample collection
Baishahe Reservoir S2 114.29 23.82 Baisha River
@@ 2.4. Model and stability analysis
According to the Routh-Hurwitz criterion, the system (1) is locally asymptotically stable at the coexistent equilibrium $E$ if the following conditions hold simultaneously: $a_{1}>0,\quad a_{2}>0,\quad a_{0}>0\quad\text{and}\quad a_{1}a_{2}>a_{0}$. If any of these conditions fails, the equilibrium $E$ is unstable.
We demonstrate this stability with a representative numerical simulation. Using the parameter set (see caption of Fig. 2 for values), the system converges to the co-existent equilibrium E(3.84, 0.137, 0.0193)E(3.84,\,0.137,\,0.0193). The eigenvalues of $\mathbf{J}(E)$ are all negative real numbers ($\lambda_{1}=-0.742$, $\lambda_{2}=-0.132$, $\lambda_{3}=-1.98\times 10^{-6}$), satisfying the stability condition. This convergence is visualized in Fig. 2, where state trajectories from different initial conditions all approach the stable equilibrium point in both time series and phase space.
@@ 3.1. Spatiotemporal dynamics of water quality parameters
Figure 3: Interannual variations (2020-2024) of (a) water temperature, (b) total nitrogen (TN), (c) total phosphorus (TP), and (d) Chl-a concentrations in Tiantangshan (S1), Baisha River (S2), and Meizhou (S3) reservoirs. Note the distinct seasonal temperature cycle, the high variability in TN, and the overall increasing trend in Chl-a.
@@ 3.3. Quantitative analysis of chlorophyll-a concentration in relation to water temperature
The simulated response of chlorophyll-a (Chl-a) concentration to temperature across the three reservoirs is presented in Fig. 6. The model predicts a unimodal relationship, with Chl-a increasing to an optimum before declining at higher temperatures. Consistent with literature values (Gobler, 2020; Paerl and Huisman, 2009), the optimal temperature range for Chl-a accumulation in our systems was approximately 20-30 ℃, with peak synthesis efficiency observed between 25-28 ℃. Notably, our simulations indicated sustained Chl-a growth across a broader range of 18-32 ℃for reservoirs S1 and S2. Concentrations in S1 and S2 were generally higher and less sensitive to temperature variation compared to S3, where Chl-a levels showed a more pronounced response to temperature fluctuations.
@@ 4. Discussion
This work has several uncertainties. (1) The differing sampling frequencies limit fine-scale comparison between reservoirs. (2) The PCA and model explain a substantial portion but not all of the observed variance, indicating missing drivers such as light availability, specific nutrient fractions, or unmeasured biotic interactions. (3) The model parameters are site-specific; transferability to other reservoirs requires further validation. (4) Future climate projections (e.g., increased temperature and rainfall variability) were not simulated, yet these factors are critical for long-term forecasting (Gobler, 2020; Fong et al., 2025). Future work should incorporate high-frequency sensor data, include biotic interactions and vertical processes, and use the model to forecast ecosystem states under climate change scenarios. Building on frameworks like regional hydro-ecological simulation system, which integrates carbon, water, and nutrient fluxes at the watershed scale, future iterations of our model could expand to a spatially distributed framework to better account for landscape heterogeneity and subsurface processes (Tague and Band., 2004). Additionally, integrating satellite-derived Chl-a data, as successfully demonstrated in other subtropical reservoirs, could provide cost-effective, high-resolution spatial validation and upscaling potential for our calibrated model, bridging the gap between point measurements and whole-reservoir management (Nthunya et al., 2018). Ultimately, such integrated approaches are essential for developing sustainable water management strategies under increasing anthropogenic and climatic pressures (Qi et al., 2025).
@@ 5. Conclusions
This study integrated five years of monitoring data with a dynamic hydro-ecological model to investigate Chl-a dynamics in three subtropical reservoirs. The key findings are:
  • Chl-a concentrations exhibited a significant increasing trend from 2020 to 2024, indicating progressing eutrophication. This trend reflects similar eutrophication challenges in subtropical freshwater systems across Asia and globally (Haq and Muhammad., 2023; Muhammad., 2023; Tokatlı et al., 2025, 2023; Muhammad et al., 2024; Mpakairi et al., 2024).
  • Multivariate analysis revealed that TN, TP, and temperature were the primary co-varying drivers, with TN showing a particularly strong association with Chl-a trends. This underscores the importance of considering nitrogen management in similar subtropical systems.
  • The developed dynamic model successfully simulated the non-linear interactions among these factors and reproduced the observed Chl-a patterns with good accuracy (R2 > 0.85), demonstrating the utility of process-based approaches.
  • The model suggests that in these specific reservoirs, managing nitrogen loads may be as crucial as controlling phosphorus to mitigate algal growth, highlighting the need for dual-nutrient strategies.
While the model is a simplification of reality and does not capture all episodic events or spatial heterogeneities, it provides a mechanistic, quantitative framework for understanding and predicting Chl-a dynamics. This work supports the move towards multi-factor, process-based modeling for the sustainable management of subtropical reservoir ecosystems in an era of global change (Maeda et al., 2019; Paerl and Huisman, 2009). Furthermore, the model was calibrated and validated on the same set of reservoirs. Its generalizability to other subtropical reservoirs remains to be tested through external validation with independent datasets.
@@ Acknowledgments
J.K. conceived and designed the project. All authors supervised the study, developed the model, analyzed the data, carried out the theoretical analysis and numerical simulations, and wrote the paper.

### v3

@@ Abstract
Eutrophication and harmful algal blooms, driven by complex interactions among nutrients and climate, threaten freshwater ecosystems globally, particularly in densely populated Asian regions where rapid urbanization and agricultural intensification exacerbate nutrient pollution. Understanding the non-linear interactions among water temperature, nutrient levels, and chlorophyll-a (Chl-a) dynamics is crucial for addressing eutrophication in freshwater ecosystems. Many existing studies, however, tend to oversimplify these relationships and lack validation with long-term field data. Here, we conducted multi-year field monitoring (2020-2024) of key environmental factors, including total nitrogen (TN), total phosphorus (TP), water temperature, and Chl-a, across three reservoirs in Guangdong Province, China: Tiantangshan (S1), Baisha River (S2), and Meizhou (S3). Chl-a concentrations showed significant spatiotemporal variability, ranging from 1.2 to 11.8 $\mu\text{g}/\text{L}$, with a general increasing trend indicative of progressing eutrophication. Strong positive correlations were found between Chl-a and TN, TP, and temperature. Numerical analysis of the long-term data revealed TN as a more influential driver than TP for Chl-a proliferation in these systems, with Chl-a increasing by an average of 4.2 $\mu\text{g}/\text{L}$ per unit increase in TN (mg/L), compared to 2.8 $\mu\text{g}/\text{L}$ per unit increase in TP (mg/L). Based on the collected data, we developed and calibrated a dynamic multi-factor hydro-ecological model. The model accurately reproduced the observed Chl-a patterns (R2 > 0.85), identifying synergistic effects between temperature and nutrients, particularly a 15% enhancement in Chl-a growth rate when temperature exceeded 25 ℃ concurrent with high TN. The model offers a robust theoretical basis for predicting Chl-a dynamics and supports science-informed management strategies, such as targeted nitrogen load reduction, to mitigate eutrophication in subtropical reservoirs. This study underscores the necessity of multi-factor, process-based approaches for sustainable water resource management under changing climatic and anthropogenic pressures.
@@
3Ocean College, Jiangsu University of Science and Technology, Zhenjiang 212100, China
@@ 1. Introduction
Aquatic ecosystems play a fundamental role in supporting life on Earth, with their structure and functioning shaped by complex interactions among key environmental factors, particularly water temperature, nutrient availability (e.g., total nitrogen (TN) and total phosphorus (TP)), and algal biomass, typically represented by chlorophyll-a (Chl-a). In recent decades, global climate change and intensified human activity have profoundly altered these ecosystems, driving more frequent extreme hydrological events, accelerating eutrophication, and contributing to widespread biodiversity loss (Fong et al., 2025; Priya et al.,, 2023; Hader and Barnes, 2019; Menden-Deuer et al., 2025; Soanar et al., 2024; Zhao et al., 2025; Hirata et al., 2025; Boyacioglu et al., 2024). In Asia, rapid urbanization and agricultural intensification have exacerbated nutrient pollution in freshwater bodies. Studies from South Asia, including Pakistan and Bangladesh, document severe eutrophication pressures in river basins and lakes, highlighting a regional crisis that demands locally calibrated understanding and solutions (Haq and Muhammad., 2023; Muhammad., 2023; Tokatlı et al., 2025, 2023; Muhammad et al., 2024). Specifically, in China’s subtropical regions, reservoirs are experiencing increasing frequency and intensity of algal blooms due to combined pressures of excessive nutrient loading from watersheds and rising water temperatures associated with climate change. Effectively managing these systems requires a quantitative understanding of how these multiple drivers interact to control phytoplankton biomass.
@@
Although multi-factor modeling in aquatic ecology has advanced in recent years, many models remain limited in predictive and explanatory power. Early studies typically focused on linear, single-factor relationships, such as nutrient concentrations and Chl-a growth. However, the emerging field of complexity science has prompted a shift toward integrative approaches that capture interactions among multiple drivers. For instance, Liu et al. (Liu et al., 2021) employed a coupled hydrodynamic-ecological model to quantify the interactive effects of TN, TP, total suspended solids (TSS), and light availability on phytoplankton competition in Chagan Lake, identifying TSS as a primary driver of cyanobacterial succession. Similarly, Qian et al. (Qian et al., 2024) applied deep learning to model algal bloom dynamics in Taihu Lake, revealing synergistic effects between thermal stratification and nutrient input. Zhang et. al. (Zhang et al., 2024) explored spatial patterns of TP and Chl-a in reservoirs, highlighting the influence of geographic and meteorological factors such as latitude, slope, and temperature on Chl-a variability. Remote sensing studies in subtropical reservoirs, such as the Nandoni reservoir in South Africa, have further corroborated the utility of high-resolution monitoring in capturing the spatial heterogeneity of Chl-a, often revealing higher concentrations near inflows and reservoir edges, which aligns with the need for spatially explicit understanding of eutrophication drivers (Mpakairi et al., 2024). Systematic reviews emphasize the growing effort to develop integrated models that couple watershed processes with in-lake dynamics to support management (Shi et al., 2024; Buta et al., 2023). In South Asia, research in the Ganges-Brahmaputra and Indus River basins has also highlighted the critical role of anthropogenic nutrient loads and climate variability in driving aquatic ecosystem changes (Muhammad et al., 2024; Tokatlı et al., 2023).
@@
While these studies contribute valuable data and demonstrate the potential of advanced modeling techniques, they often rely heavily on statistical or machine learning approaches, offering limited insight into the underlying ecological mechanisms  (Liu et al., 2021; Qian et al., 2024; Zhang et al., 2024). A key limitation of many existing models lies in their calibration with short-term datasets and the consequent lack of rigorous validation against long-term observational records; this deficiency critically undermines their ability to accurately simulate and predict system responses under realistic conditions of interannual climatic variability and episodic disturbance events—such as the extreme rainfall prevalent in subtropical monsoonal regions—which can drastically alter nutrient fluxes and phytoplankton dynamics (Buta et al., 2023). Moreover, traditional models such as the Water Erosion Prediction Project (WEPP), although effective in simulating physical processes like soil erosion, fall short in capturing hydro-ecological interactions (Flanagan et al., 2012; Chen et al., 2025). Some progress has been made by integrating robust optimization and scenario analysis into water resource management, for example, Xu et al. (Xu et al., 2022) incorporated multi-objective optimization and copula-based uncertainty quantification into cascade reservoir planning, but the ecological models remain relatively simplified. In contrast, emerging dynamic multi-factor models offer promising tools to explore non-linear feedbacks and temporal continuity among interacting ecological variables, including Chl-a growth, nutrient cycling, and temperature dynamics. Despite their potential, these models are still in early stages of development and are seldom applied to real-world aquatic systems.
@@
To address these issues, this study aims to: (1) quantify the spatiotemporal dynamics and key drivers of Chl-a in three subtropical reservoirs using a comprehensive five-year (2020-2024) monitoring dataset; (2) develop a process-based, dynamic multi-factor model that mechanistically integrates the effects of water temperature, TN, and TP on Chl-a; and (3) calibrate and validate the model against the long-term data to assess its predictive capability and provide a tool for exploring management scenarios. We propose a dynamic, multi-factor modeling approach to investigate the interactions among water temperature, TN, TP, and Chl-a in three freshwater reservoirs in Guangdong, China. By leveraging long-term monitoring data and integrating ecological mechanisms into the modeling framework, this work aims to deepen our understanding of eutrophication processes and provide a scientific basis for informed aquatic ecosystem management.
@@ 2.1. Research sites and sample collection
Baisha River Reservoir S2 114.29 23.82 Baisha River
@@ 2.4. Model and stability analysis
According to the Routh-Hurwitz criterion, the system (1) is locally asymptotically stable at the coexistent equilibrium $E$ if the following conditions hold simultaneously: $a_{1}>0,\quad a_{2}>0,\quad a_{0}>0\quad\text{and}\quad a_{1}a_{2}>a_{0}$. If any of these conditions fails, the equilibrium $E$ is unstable. Using the parameter set (as specified in the caption of Fig. 2), the system converges to a stable fixed point EE(see Fig. 2). At this equilibrium, all eigenvalues of the Jacobian matrix $\mathbf{J}(E)$ are negative real parts, satisfying the condition for local asymptotic stability. This convergence is visualized in Fig. 2, where trajectories from different initial conditions approach the stable fixed point in both time series and phase space.
@@ 3.1. Spatiotemporal dynamics of water quality parameters
Figure 3: Interannual variations (2020-2024) of total nitrogen (TN), total phosphorus (TP), and Chl-a concentrations in Tiantangshan (S1), Baisha River (S2), and Meizhou (S3) reservoirs. Note the distinct seasonal temperature cycle, the high variability in TN, and the overall increasing trend in Chl-a.
@@ 3.3. Quantitative analysis of chlorophyll-a concentration in relation to water temperature
The simulated response of Chl-a concentration to temperature across the three reservoirs is presented in Fig. 6. The model predicts a unimodal relationship, with Chl-a increasing to an optimum before declining at higher temperatures. Consistent with literature values (Gobler, 2020; Paerl and Huisman, 2009), the optimal temperature range for Chl-a accumulation in our systems was approximately 20-30 ℃, with peak synthesis efficiency observed between 25-28 ℃. Notably, our simulations indicated sustained Chl-a growth across a broader range of 18-32 ℃for reservoirs S1 and S2. Concentrations in S1 and S2 were generally higher and less sensitive to temperature variation compared to S3, where Chl-a levels showed a more pronounced response to temperature fluctuations.
@@ 4. Discussion
This work acknowledges several limitations and uncertainties. (1) The differing sampling frequencies between reservoirs constrain direct fine-scale temporal comparisons. (2) Although the PCA and model captured a substantial portion of the variance, the remaining unexplained variance points to the influence of unmeasured drivers, such as light availability, specific nutrient fractions (e.g., bioavailable phosphorus), or complex biotic interactions. (3) The model parameters were calibrated for the specific conditions of the studied reservoirs; therefore, their direct applicability to other systems without local validation remains an uncertainty. (4) The current modeling framework does not incorporate projections of future climate, which is a key factor for long-term predictions given the sensitivity of algal growth and nutrient cycles to warming and hydrological changes (Gobler, 2020; Fong et al., 2025; Soanar et al., 2024). These limitations underscore the critical need for the integrated and expanded research pathways detailed in the dedicated Future Perspectives section. Advancing along these multifaceted lines is imperative to formulate resilient and sustainable management strategies for freshwater ecosystems facing escalating pressures (Qi et al., 2025).
5. Future Perspectives
This work opens several avenues for future research. (1) The differing sampling frequencies limit fine-scale comparison between reservoirs. Future monitoring should adopt harmonized, high-frequency (e.g., bi-weekly or continuous sensor-based) sampling to better resolve short-term dynamics and event responses, as emphasized in studies of long-term reservoir water quality evolution (Buta et al., 2023). (2) The PCA and model explain a substantial portion but not all of the observed variance, indicating missing drivers. Future model iterations should incorporate additional factors such as photosynthetically active radiation (PAR), silicate concentrations, zooplankton grazing pressure, and microbial loop dynamics, which are known to mediate phytoplankton community structure (Menden-Deuer et al., 2025). (3) The model parameters are site-specific; transferability to other subtropical reservoirs requires further validation with independent datasets from diverse systems. (4) Incorporating future climate projections (e.g., increased temperature and altered precipitation patterns from CMIP6 models) into the dynamic framework is critical for forecasting long-term ecosystem states and resilience, given the profound influence of climate on algal blooms and nutrient cycling (Gobler, 2020; Fong et al., 2025; Soanar et al., 2024). (5) Expanding the model to a spatially distributed framework (Tague and Band., 2004), potentially integrated with remote sensing data for surface Chl-a validation (Nthunya et al., 2018), would allow for the simulation of horizontal heterogeneity and identification of critical source areas within reservoirs. (6) Ultimately, linking this reservoir model with watershed-scale nutrient loading models (e.g., SWAT) would create a powerful integrated tool for evaluating the effectiveness of land-based management actions on downstream water quality, aligning with integrated catchment management approaches (Chen et al., 2025).
@@ 6. Conclusions
[…]
Overall, this study underscores the value of combining long-term observational data with process-based dynamic modeling to understand and predict complex eutrophication dynamics. It highlights the necessity of dual-nutrient (N and P) strategies and temperature-aware management in the sustainable governance of freshwater resources in rapidly developing subtropical regions under climate change.
@@ Acknowledgments
Haizhao Guan: Writing – review & editing, Writing – original draft, Visualization, Methodology, Investigation, Formal analysis, Data curation, Conceptualization. Yiyuan Niu: Writing – review & editing, Writing – original draft, Visualization, Methodology, Investigation, Formal analysis. Chuanjin Zu: Writing – review & editing, Investigation, Funding acquisition. Ju Kang: Writing – review & editing, Writing – original draft, Visualization, Validation, Software, Methodology, Investigation, Formal analysis, Data curation, Conceptualization.